### v1

@@ 2.7 Gas accretion by individual stars
where $j$ is the minimum integer for which the condition $dt_{\rm new}<dt_{\rm req}$ is fulfilled. On the other hand, if $dt_{\rm req}\geq 2dt$, then $dt_{\rm new}=2dt$.
@@ 3.1 Cluster and core contraction
with $N$ being the number of stars and $t_{\rm cross}$ the crossing time. We therefore expect $t_{\rm rh}$ to be longer in a cluster with more stars (provided $t_{\rm cross}$ remains unchanged). Thus, when $\epsilon_{\rm SF}$ is higher, $N$ is larger and $t_{\rm rh}$ is longer. We also find that the contraction rate is slower in clusters with rapid SF. The same argument applies here; the number of stars during the time span considered for this analysis is higher, which increases the relaxation time. As a result, the cluster contracts more slowly.
@@ 3.3 Emergence of the most massive object
Our simulations show that most of the stars in the high-mass end of the IMF, i.e.,with an initial mass $M>100$ M, typically form outside the core and therefore never experience a collision. In fact, the fraction of these stars that survive until the end of the simulation is typically $>98$%. Moreover, stars that experience a collision are typically the ones that are formed in between the cluster core and the half-mass radius, whereas the majority of the stars formed in the core ($\gtrsim 90$%) survive until the end of the simulation. This means that even though the collisions occur in the cluster core, the colliding stars are typically formed outside.
This is somewhat different in model E6, in which the clusters experience core collapse. In this case, the survival rate of $M>100$ M stars is $\sim 95\%$, and the majority of the high mass stars in the core ($>60$%) experience a collision.
@@ 4 Conclusion and Discussion
Our models of Pop III star clusters are based on realistic initial conditions taken directly from a cosmological simulation. We have developed a model for their subsequent assembly and evolution. We find that BHs of up to a few thousand solar masses are formed through this scenario. These BH seeds are heavier compared to the single Pop III remnant scenario. Additionally, the fact that the BH is embedded in a massive star cluster will further aid its sinking into the deep potential well of the galaxy (Mukherjee et al. 2024). In particular, we might expect the system of stars and BH to experience subsequent episodes of gas accretion. BHs in such environments have been shown to grow at rates that may exceed the Eddington limit, with e-folding times of 200-300 Myr (Partmann et al. 2025). This pathway, therefore, offers a promising alternative to the DCBH scenario. The higher abundance of BH seeds forming through this channel can explain the recently observed abundance of AGN at high $z$ as found in high-resolution cosmological simulations (Mayer et al. 2025), along with the potential for further efficient growth. However, certain effects are still not included in our models, and we discuss some of these caveats below.
We have neglected mass loss due to stellar winds, given the zero metallicity of the stars considered in our models (Krtička & Kubát 2006). It is worth noting that although rotating Pop III star models experience mass loss through winds, this only occurs once the stars become red giants (Meynet et al. 2006). Therefore, our results are robust against this effect, particularly with respect to the final masses of black holes, which would form via direct collapse (Heger et al. 2003).

### v2

@@ 2.7 Gas accretion by individual stars
where $j$ is the minimum integer for which the condition $dt_{\rm new}<dt_{\rm req}$ is fulfilled. On the other hand, if $dt_{\rm req}\geq 2dt$, then $dt_{\rm new}=2dt$. We note that this timestep reduction is rarely triggered in our simulations and is only relevant for high-mass stars.
@@ 3.1 Cluster and core contraction
with $N$ being the number of stars and $t_{\rm cross}$ the crossing time. We therefore expect $t_{\rm rh}$ to be longer in a cluster with more stars (provided $t_{\rm cross}$ remains unchanged). Thus, when $\epsilon_{\rm SF}$ is higher, $N$ is larger and $t_{\rm rh}$ is longer. We also find that the contraction rate between 0.25-0.6 Myr is slower in clusters with rapid SF. The same argument applies here; the number of stars during the time span considered for this analysis is higher, which increases the relaxation time. As a result, the cluster contracts more slowly.
@@ 3.3 Emergence of the most massive object
Our simulations show that most of the stars in the high-mass end of the IMF, i.e.,with an initial mass $M>100$ M, typically form outside the core and therefore never experience a collision. In fact, the fraction of these stars that survive until the end of the simulation is typically $>98$%. Moreover, stars that experience a collision are typically the ones that are formed in between the cluster core and the half-mass radius, whereas the majority of the stars formed in the core ($\gtrsim 90$%) survive until the end of the simulation. This means that even though the collisions occur in the cluster core, the colliding stars are typically formed outside. This is somewhat different in model E6, in which the clusters experience core collapse. In this case, the survival rate of $M>100$ M stars is $\sim 95\%$, and the majority of the high mass stars in the core ($>60$%) experience a collision.
@@ 4 Conclusion and Discussion
Our models of Pop III star clusters are based on realistic initial conditions taken directly from a cosmological simulation. We have developed a model for their subsequent assembly and evolution. We find that BHs of up to a few thousand solar masses are formed through this scenario. These BH seeds are heavier compared to the single Pop III remnant scenario. Additionally, the fact that the BH is embedded in a massive star cluster will further aid its sinking into the deep potential well of the galaxy (Mukherjee et al. 2025). In particular, we might expect the system of stars and BH to experience subsequent episodes of gas accretion. BHs in such environments have been shown to grow at rates that may exceed the Eddington limit, with e-folding times of 200-300 Myr (Partmann et al. 2025). This pathway, therefore, offers a promising alternative to the DCBH scenario. The higher abundance of BH seeds forming through this channel can explain the recently observed abundance of AGN at high $z$ as found in high-resolution cosmological simulations (Mayer et al. 2025), along with the potential for further efficient growth. However, certain effects are still not included in our models, and we discuss some of these caveats below.
We have neglected mass loss due to stellar winds, given the zero metallicity of the stars considered in our models (Krtička & Kubát 2006). It is worth noting that although rotating Pop III star models experience mass loss through winds, this only occurs once the stars become red giants (Meynet et al. 2006). Therefore, our results are robust against this effect, particularly with respect to the final masses of black holes, which would form via direct collapse (Heger et al. 2003). We did not include the effect of dynamical friction in our models presented here. Doing this will lead to a faster sinking of massive stars into the cluster core and subsequently may increase the total number of collisions. This may ultimately increase the mass of the MMO.
We have also neglected the stellar feedback during AMUSE simulations which could increase the gas temperature and prevent further star formation. However, previous studies show that the SN feedback is unable to expel the gas from massive embedded star clusters (Krause et al. 2016; Grudić et al. 2018; Lahén et al. 2025). Moreover, star formation can still continue in such system despite pre-SN feedback as the gas forms a disc in the cluster core (Lahén et al. 2025). Therefore, we expect that our results would not change significantly by such effects.
@@ Appendix A Orbital parameters that led to stellar collisions
Appendix B Gas inflow and halo growth
The gaseous collapse in the minihalo was prevented in a similar fashion to Latif et al. (2022b) and Cold dense gas inflows resulted in rapid gas accretion as shown in Fig. 11. Turbulent motions prevented gas infall until the halo grew up to a few times $10^{7}$ M, and large inflow velocities overcame the turbulent motions leading to catastrophic collapse, see Fig. 12. We also present a snapshot of the halo on a scale of the virial radius in Fig. 13
Figure 11: Large scale view of the halo showing its environment and cold dense streams of gas that drive turbulent motions at the center.
Figure 12: Gas velocity profiles are shown against the enclosed mass. Dotted lines show the turbulent velocities, solid lines the infall velocities and dashed lines thermal sound speeds. Magenta, green and red colors show time evolution (from earlier to later times), respectively.
Figure 13: View of the halo on a scale corresponding to the virial radius, which is equal to 260 pc.